# Advances in Kidney Biopsy Lesion Assessment through Dense Instance Segmentation

### v2

@@ 4 Results
In this section, we start with an overview of the kidney biopsy dataset used in this study. We then proceed to draw a fair comparison between DiffRegFormer against established end-to-end instance segmentation models specifically designed to handle dense, multi-class, and multi-scale objects at the ROI level. To ensure consistency, all benchmark methods are reproduced utilizing the mmdetection package ([7]). Further, we evaluate the performance of our lesion classifier on cropped images of anatomical structures, derived from the predicted instance masks by DiffRegFormer. Furthermore, we conduct a complexity analysis of DiffRegFormer to identify an optimal balance between its performance and computational efficiency. Finally, we present extensive ablation studies on DiffRegFormer to evaluate the impact of various training strategy combinations.
@@ 4.1 Datasets
Biopsy specimens were prepared in accordance with the Pathology Laboratory Protocol for kidney biopsies. Tissues were collected via a standard procedure of operations: core needle biopsy, formalin-fixed, rapidly processed by a routine tissue processor, embedded in paraffin, serially sectioned at $3-4\mu m$, mounted onto adhesive slides, and stained with Jones silver staining. The dataset comprises $148$ patient biopsy samples (both transplant and native kidneys) collected from the multicenter’s archive of the Departments of Pathology at LUMC, AUMC, UMCU in the Netherlands or the archive of the Department of Nephrology at Leuven in Belgium. Prior to analysis, all biopsies were anonymized by a pathology staff member. Annotation was performed employing ASAP version $1.9$ for Windows [^5] https://computationalpathologygroup.github.io/ASAP/ and our custom software Slidescape [^6] https://github.com/amspath/slidescape. The $148$ Jones-stained WSIs were acquired from scanners at a resolution of $\mathrm{PPM}=0.25$ in BIG-TIFF format [^7] https://www.awaresystems.be/imaging/tiff/bigtiff.html. Openslide [^8] https://openslide.org/ was used to extract $303$ ROIs at level 00 of the BIG-TIFF images. Table.3 outlines the dataset composition, and Figure.8 illustrates the distribution of ROI short sizes. Additionally, $115$ PAS-stained WSIs from LUMC were utilized solely for testing of domain transferability. Notably, ROIs from the same WSIs were exclusively allocated to either training or validation. To train the lesion classifier, $21,889$ patches of individual objects were extracted from $303$ ROIs, including $916$ glomeruli patches with binary lesion labels ($382$ with sclerotic and $534$ healthy) and $20,973$ tubuli patches with binary labels ($10,647$ with atrophy and $10,326$ healthy), as detailed in Figure.6.
Table 3: Instance Data Set Configuration.
Source Scanner WSI ROIs Train Eval
LUMC Philips UFS 65 142 130 12
@@
Fig. 8: The histogram of the shortest edges of ROIs extracted from WSIs. In our dataset, the shortest length is 387 and the largest one is 4565.
Fig. 9: A visualization for all methods with Jones-stained kidney biopsies. Our model can detect objects of various sizes and shapes. Over each region, decent instance masks can be generated although there are some overlays between different instances. Mask-RCNN-based models cannot locate one large object within one box because they do not adopt attention mechanisms to capture long-range dependencies. QueryInst with global queries cannot process dense objects since they fail to detect every object within one biopsy.
Fig. 10: A visualization for our model on the PAS-stained kidney biopsies with increasing degrees of fibrosis, inflammation (left to right), and variation in stain intensity.
Fig. 11: Complexity Analysis on Jones-stained datasets.
@@
Fig. 12: Ablation study on combinations of different strategies.
Table 4: Evaluation on combinations of different strategies
@@
  • 1.
    glo: glomeruli. art: arteries. tub: tubuli. all: overall. det denotes detection. ins means instance. ”-” means not working.
  • 2.
    feature: if use seperate feature maps between bbox-decoder and mask-decoder. boxes: if use ground-truth boxes in the sampling process. sample: if maintain class balance in the sampling process.
  • 3.
    The symbol ✓ denotes taking one strategy while ✗ means taking the converse strategy.
Table 5: Evaluation of detection performances on different types of queries
# 200 # 300 # 400 # 500 # 600
@@
atrophy normal sclerosis normal
train 10647 10326 382 534
eval 568 906 22 51
@@
Fig. 13: A visualization of ablation study on query types.
Fig. 14: A visualization of DiffRegFormer predicts dense objects on one Whole Slide Image (WSI).
Fig. 15: A visualization of lesion classifier predicts the probability of class-wise lesion on ROIs.
@@ 4.2 Experiments
In our experimental analysis, we conducted a fair comparison of our proposed DiffRegFormer against well-established end-to-end instance segmentation models: Mask R-CNN ([21]), Cascade Mask R-CNN ([4]), and QueryInst ([17]), specifically designed to simultaneously handle dense, multi-class, multi-scale objects at the ROI level. All the models under comparison employed a ResNet-$50$ backbone pre-trained on the ImageNet ([22]) and were trained on $249$ ROIs over $40000$ iterations. We utilized the mmdetection data augmentation pipeline: each ROI was first subjected to RandomFlip with a $0.5$ probability, followed by random selection between RandomResize or RandomCrop also with a probability $0.5$. RandomResize adjusted the ROI to a size with the shortest edge between $480$ and $1333$ pixels, while RandomCrop extracted patches with the shortest edge between $384$ and $600$ pixels. Finally, pixel values were normalized to $(-1,1)$. As mmdetection supports training with variable image sizes, all models were allowed to train on objects at multiple scales.
@@
For feature extraction, we used $256$ channels for queries. During both training and inference, DiffRegFormer employed 500 dynamic queries, while QueryInst used $300$ and $500$ global queries respectively. Cascade Mask-RCNN, QueryInst, and our method utilized $6$ cascading stages. Table.1 presents performance comparisons for object detection w.r.t bounding boxes and instance segmentation w.r.t masks respectively. For all experiments, DiffRegFormer did not utilize refinement during inference (i.e., iteration set to 1). Notably, Mask-RCNN ([21]) and cascade Mask-RCNN ([4]) are two-stage methods employing RPN networks instead of dynamic queries. QueryInst is a one-stage method with global queries. In contrast, DiffRegFormer is categorized as a one-stage method with dynamic queries.
@@
Detection. DiffRegFormer achieves a detection Average Precision (AP) of 52.1% with a ResNet-50 backbone, surpassing competitors such as QueryInst, Mask R-CNN, and Cascade Mask R-CNN by notable margins. Intriguingly, iterative refinement appears to hinder detection performance for small objects while benefiting the detection of medium to large objects. This suggests the need for additional operations to mitigate information loss for small objects during the refinement process. Furthermore, the use of global queries appears less effective for handling dense objects, particularly when the number of queries significantly exceeds the number of feature channels ($300$ or $500$ compared to $256$), leading to a marked decline in detection performance.
@@
Instance segmentation. DiffRegFormer achieves a 46.8% AP for instance segmentation, outperforming its counterparts. Likewise, the application of iterative refinement to bounding boxes is found to significantly impact the instance segmentation outcomes. This effect can be attributed to the binary segmentation process within the boxes, where the quality of the instance masks is inherently linked to the recall and precision rates of the bounding boxes.
@@
Table.2 presents the performance per object class. Our model demonstrates an overall superior ability to detect all anatomical structures compared to other methods. Additionally, it excels in the instance segmentation of glomeruli and arteries by notable margins and shows a slight advantage in tubuli segmentation. These findings indicate our model’s ability to process dense, multi-class objects with diverse scales at the ROI level. However, we observe a performance bottleneck for all models in the processing of arteries, which poses a limitation for clinical application due to the significant size variation within the dataset. For instance, the largest arteries are hundreds of times larger than the smallest. This fact suggests the need for more sophisticated attention mechanisms to capture global contextual information for such extreme scale variations.
@@
Figure.9 demonstrates DiffRegFormer’s capability to effectively detect objects at various scales. It accurately generates instance masks within bounding boxes, demonstrating precise localization of larger arteries (see the second row of the figure). In contrast, Cascade Mask R-CNN has difficulty in locating larger objects within a single bounding box, due to the lack of an attention mechanism for capturing long-range dependencies. Furthermore, Mask-RCNN tends to generate multiple detections for a single object(as seen in the third row), resulting from the absence of cascade refinement. These observations suggest that iterative bounding box refinement during training significantly reduces multiple predictions for a single object, while the attention mechanism with dynamic queries processes large-scale objects. Nevertheless, QueryInst, with its reliance on global queries, may struggle to detect all instances effectively in datasets with significant size and shape variations. QueryInst-$300$ tends to miss many structures, while QueryInst-$500$ can even degenerate into generating multiple false predictions for a single object. Furthermore, RCNN-based models share a limitation: a single pixel may be allocated to multiple instance masks due to pixel-wise predictions within each independent ROI. This inherently prevents enforcing the spatial exclusivity of pixels to instances. Additionally, Figure.10 demonstrates DiffRegFormer’s potential for domain transfer, as it generates reasonable results on PAS-stained images despite being trained exclusively on Jones’ silver-stained images. Finally, Figure.14 highlights that DiffRegFormer, when combined with techniques such as sliding window and stitching, can be effectively applied to WSIs, further demonstrating its effectiveness.
@@
Lesion classification. The data augmentation pipeline for training the lesion classifier is relatively straightforward. We set the batch size as $96$ for the input crops, where $32$ crops for each class. Each crop first undergoes RandomFlip with a probability of $0.5$, followed by resizing to $256\times 256$. Finally, pixel values are normalized to the range of $(-1,1)$. As shown in Table.6, the classifier achieves an overall precision of $89.2\%$ precision and a recall of $64.6\%$. More specifically, per-class precision for tubuli atrophy is $85.1\%$, and $93.2\%$ for glomeruli sclerosis. In addition, per-class recall is $40.5\%$ for tubuli atrophy and $88.7\%$ for glomeruli sclerosis. Since the lesion classifier operates on crops of anatomical structure generated by DiffRegFormer during inference, its recall rate is strongly influenced by the performance of our dense instance segmentation model. In contrast, the high precision indicates that lesion classification on correctly cropped structures is highly accurate. In summary, enhancing the performance of DiffRegFormer is expected to further improve the lesion classifier’s recall rate. Figure.15 visualizes the mapping of lesion predictions onto the instance segmentation results.
4.3 Complexity Analysis

### v3

@@ 4 Results
This section begins with an overview of the kidney biopsy dataset used in this study. We then make a fair comparison between DiffRegFormer and established end-to-end instance segmentation models specifically designed to handle dense, multi-class, and multi-scale objects at the ROI level. In addition, in order to include retrospective reflection in our analysis, we have compared our model with a semantic segmentation based method [28]. The results of this comparison are presented in section 4.2.2 and the tables therein. All benchmark methods are reproduced utilizing the mmdetection package [7] to ensure consistency. Further, we evaluate the performance of our lesion classifier on cropped images of anatomical structures derived from the instance masks predicted by DiffRegFormer. Additionally, we have conducted a complexity analysis of DiffRegFormer to identify an optimal balance between its performance and computational burden. Finally, we present extensive ablation studies on DiffRegFormer that evaluate the impact of various training strategies.
@@ 4.1 Datasets
The biopsy specimens were prepared according to the Pathology Laboratory Protocol for kidney biopsies. The tissues were collected by a standard procedure of operations: core needle biopsy, formalin fixed, rapidly processed by a routine tissue processor, embedded in paraffin, serially sectioned at $2-4\mu m$, mounted onto adhesive slides, and stained with Jones silver staining. The dataset comprises $148$ patient biopsy samples (both transplant and native kidneys) collected from the multi-center archive of the Departments of Pathology at LUMC, AUMC, UMCU in the Netherlands or the archive of the Department of Nephrology at Leuven in Belgium. Before analysis, all biopsies were anonymized by a pathology staff member. Annotation was performed using the ASAP version $1.9$ for Windows [^5] https://computationalpathologygroup.github.io/ASAP/ and our custom software Slidescape [^6] https://github.com/amspath/slidescape. The $148$ Jones-stained renal WSIs were acquired from scanners at a resolution of $\mathrm{PPM}=2$ in BIG-TIFF format [^7] https://www.awaresystems.be/imaging/tiff/bigtiff.html. Openslide [^8] https://openslide.org/ was used to extract $303$ ROIs at level 00 of the BIG-TIFF images. Table 3 outlines the dataset composition, and Figure 8 illustrates the distribution of ROI short sizes. Furthermore, LUMC $115$ PAS-stained renal WSIs were used only to test stain transfer. In particular, ROIs from the same WSIs were exclusively allocated to either training or validation. To train the lesion classifier, $21,889$ patches of individual objects were extracted from $303$ ROIs, including $916$ glomeruli patches with binary lesion labels ($382$ with sclerosis and $534$ normal) and $20,973$ tubuli patches with binary labels ($10,647$ with atrophy and $10,326$ normal), as detailed in Figure 6.
Table 3: Origin of datasets digitized with different scanners
Source Scanner WSIs ROIs Train Eval
LUMC [^9] Leiden University Medical Center Philips UFS 65 142 130 12
@@
Fig. 8: The histogram of the shortest edges of ROIs extracted from WSIs. In our dataset, the shortest ROI length is 387 pixels, and the largest is 4565 pixels.
Fig. 9: A visual comparison of different instance segment methods on Jones-stained kidney biopsies. It illustrates that our model, i.e., DiffRegFormer, outperforms in detecting objects of diverse size and shape. Despite some instances overlapping, it generates precise instance masks over each region. Mask-RCNN-based models struggle to localize one large object within a single bounding box as they lack attention mechanisms to capture long-range dependencies. QueryInst with static queries fails to process dense objects since it fails to detect objects of each class within one ROI.
Fig. 10: A visualization of results of our model (DiffRegFormer) on the PAS-stained kidney biopsies with increasing degrees of fibrosis and inflammation (from left to right), along with variations in stain intensity.
Fig. 11: Complexity Analysis on Jones-stained renal biopsy dataset. Figure (a) is the sample occupation across our dataset. Figure (b) is a violin plot showing the sample number distribution per class. Figure (c) shows trade-offs between performance and efficiency for different experiment settings. The red square is the optimal setting, while the green circles are the sub-optimal settings.
@@
Fig. 12: Ablation study on combinations of different strategies. The images on the left are ground-truth. The images on the right are our predictions.
Table 4: Evaluation on combinations of different strategies. The symbol ✓ denotes taking one strategy while symbol ✗ means taking the converse strategy. The best results are highlighted in bold.
glo: glomeruli. art: arteries. tub: tubuli. all: overall. det: detection. ins: instance. ”-”: not working.
feature: if use separate feature maps between bbox-decoder and mask-decoder.
boxes: if use ground-truth boxes in the sampling process.
sample: if have class balance in the sampling process.
@@
Table 5: Evaluation of detection performances on different types of queries. $\ast$ best result of each type is highlighted in bold.
mAP number 200 300 400 500 600
@@
normal atrophy normal sclerosis
train 10326 10647 534 382
eval 906 568 51 22
@@
Fig. 13: A visual comparison from an ablation study on different query types. Dynamic queries outperform static queries on detection. It should be noted that the static queries fail to detect objects of all classes within one ROI.
Fig. 14: A representative visualization of results from DiffRegFormer stitches showing predicted dense instance objects from a WSI of a Jones-stained renal biopsy.
Fig. 15: A visualization of results of the lesion classifier predicting the probability of class-wise lesions on an ROI. Combined with dense instance segmentation, our model can generate per-instance lesion identification.
@@ 4.2 Experiments
In our experimental analysis, we conducted a fair comparison of our proposed DiffRegFormer against well-established end-to-end instance segmentation models: Mask R-CNN [26], Cascade Mask R-CNN [4], and QueryInst [20], specifically designed to simultaneously handle dense, multi-class, multi-scale objects at the ROI level. All the models under comparison employed a ResNet-$50$ [27] backbone pre-trained on the ImageNet [15] and were trained on $249$ ROIs over $40000$ iterations. We utilized the mmdetection data augmentation pipeline: each ROI was first subjected to RandomFlip with a $0.5$ probability, followed by random selection between RandomResize or RandomCrop also with a probability $0.5$. RandomResize adjusted the ROI to a size with the shortest edge between $480$ and $1333$ pixels, while RandomCrop extracted patches with the shortest edge between $384$ and $600$ pixels. Finally, pixel values were normalized to $(-1,1)$. As mmdetection supports training with variable image sizes, all models were allowed to train on objects at multiple scales.
@@
For feature extraction, we used $256$ channels for queries. During both training and inference, DiffRegFormer employed 500 dynamic queries, while QueryInst used $300$ and $500$ static queries, respectively. Cascade Mask-RCNN, QueryInst, and our method utilized $6$ cascading stages. Table 1 presents performance comparisons for object detection w.r.t bounding boxes and instance segmentation w.r.t masks, respectively. For all experiments, DiffRegFormer did not utilize refinement during inference (i.e., iteration set to 1). Notably, Mask-RCNN [26] and cascade Mask-RCNN [4] are two-stage methods employing RPN networks instead of dynamic queries. QueryInst is a one-stage method with static queries. In contrast, DiffRegFormer is categorized as a one-stage method with dynamic queries.
@@
Detection. DiffRegFormer achieves an Average Precision (AP) of 52.1% for detection with a ResNet-50 backbone, thereby surpassing competitors such as QueryInst, Mask R-CNN, and Cascade Mask R-CNN by significant margins. We observe that the iterative refinement hinders detection performance for small objects while benefiting the detection of medium to large objects. It suggests additional operations to mitigate information loss for small objects during the refinement process. Furthermore, static queries are less effective for handling dense objects, particularly when the number of queries significantly exceeds the number of feature channels ($300$ or $500$ compared to $256$), leading to a marked decline in detection performance.
@@
Instance segmentation. DiffRegFormer achieves a 46.8% AP for instance segmentation and thus outperforms its counterparts. Moreover, applying iterative refinement to bounding boxes significantly impacts the instance segmentation outcomes. This effect can be attributed to the binary segmentation process within the boxes, where the quality of the instance masks is inherently linked to the recall and precision rates of the bounding boxes.
@@
Table 2 presents the performance per object class. Our model demonstrates an overall superiority in detecting all anatomical structures compared to other methods. Additionally, it performs better in the instance segmentation of glomeruli and arteries while keeping a slight advantage in tubuli segmentation. These findings indicate the ability of our model to process dense, multi-class objects with diverse scales at the ROI level. However, for each model, we observe a performance bottleneck in the processing of arteries, posing a limitation for clinical application due to the significant size variation within the dataset. For instance, the largest arteries are two orders of magnitude larger than the smallest. This fact suggests that more sophisticated attention mechanisms are needed to capture global contextual information for these extreme variations in scale.
@@
Figure 11 demonstrates DiffRegFormer’s capability to detect objects at various scales effectively. It accurately generates instance masks within bounding boxes, demonstrating precise localization of larger arteries (see the second row). In contrast, Cascade Mask R-CNN has difficulty locating larger objects within a single bounding box due to the lack of an attention mechanism for capturing long-range dependencies. Furthermore, Mask-RCNN tends to generate multiple detection results for a single object(see the third row) due to the absence of cascade refinement. These observations suggest that iterative bounding box refinement during training significantly reduces multiple predictions for a single object. At the same time, the attention mechanism with dynamic queries processes large-scale objects. Nevertheless, QueryInst, with its reliance on static queries, may struggle to effectively detect all instances in datasets with significant size and shape variations. QueryInst-$300$ tends to miss many structures, while QueryInst-$500$ can even degenerate into generating multiple false predictions for a single object. Furthermore, RCNN-based models share a limitation: a single pixel may be allocated to multiple instance masks due to pixel-wise predictions within each independent ROI. It inherently prevents enforcing the spatial exclusivity of pixels to instances. Additionally, Figure 11 depicts DiffRegFormer’s potential for domain transfer, as it generates reasonable results on PAS-stained images while being trained exclusively on Jones’ silver-stained images. Moreover, Figure 14 highlights that DiffRegFormer, combined with sliding window and stitching techniques, can be effectively applied to WSIs, further demonstrating its effectiveness.
@@
Lesion classification. The data augmentation pipeline for training the lesion classifier is relatively straightforward. We set the batch size as $96$ for the input croppings, where $32$ patches for each class. Each cropping first undergoes RandomFlip with a probability of $0.5$, followed by resizing to $256\times 256$ pixels. Finally, pixel values are normalized to the $(-1,1)$ range. As shown in Table 6, the classifier achieves an overall precision of $89.2\%$ and a recall of $64.6\%$. More specifically, per-class precision for atrophic tubuli is $85.1\%$, and $93.2\%$ for sclerotic glomeruli. In addition, per-class recall is $40.5\%$ for atrophic tubuli and $88.7\%$ for sclerotic glomeruli. Since the lesion classifier operates on croppings of the anatomical structure generated by DiffRegFormer during inference, its recall rate heavily depends on the performance of our dense instance segmentation model. However, the high precision indicates that lesion classification on correctly cropped structures is very accurate. In summary, the enhancement of the performance of DiffRegFormer is expected to improve the recall of the lesion classifier further. Figure 15 visualizes the mapping of lesion predictions onto the instance segmentation results. It reveals that our model can be a solid foundation for further clinical diagnosis.
4.2.1 Metrics for Instance Segmentation
We have computed a set of commonly used metrics to assess the performance of our method in a broader scope. These are DICE, MASD (mean average surface distance), FLOPS, inference speed, and amount of parameters. In Table 7, we can observe from the values of DICE and MASD that our method outperforms other approaches on segmentation performance. The resource efficiency is, however, moderate. The main cause of resource cost is the attention mechanism, as seen in QueryInst and DiffRegFormer. To that end, it is beneficial to adopt more efficient attention algorithms to accelerate the inference speed and lessen resource burden, tailoring for the practical clinical agenda. In addition, the arteries that vary across multiple scales make it difficult to obtain high-quality instance masks. All methods have illustrated a large average surface distance between the predicted and ground-truth boundaries.
Table 7: Overall Metrics for method assessment. T stands for trillion floating-point operations per second. sec means second in time. Para refers to the number of parameters where M denotes million.
MASD DICE FLOPS (T) Speed (sec) Para (M)
one-stage QueryInst_300 7.49 42.6 0.41 1.425 173
QueryInst_500 10.95 35.9 0.55 1.647 246
two-stage Mask-RCNN 4.78 50.7 0.20 0.864 43.98
CMask-RCNN 3.87 52.1 1.72 1.156 77.03
our model DiffRegFormer 3.64 56.6 0.55 1.324 140
4.2.2 Comparison with Segmentation Based Method
In order to supplement our analysis in terms of completeness, we have done a retrospective analysis reproducing a semantic segmentation based instance segmentation framework [28]. We denote this method as a semantic method and compare its results with ours. In Table 8, an overall comparison is depicted. Our method outperforms the semantic method in prediction quality at the cost of more inference time and computational resources. In Table 9, the superiority of our method per class is depicted. Furthermore, Figure 16 illustrates that the semantic method is vulnerable to the segmentation quality of the auxiliary boundary and interstitium class, which is insufficient in separating objects tightly touching. Such touching objects are typical in kidney tissue. Our method, on the contrary, can detect and segment each object.
Table 8: Overall evaluation result. The semantic method is an instance segmentation framework that combines semantic segmentation with a post-processing operation. We have reproduced it following the literature [28]. AP denotes average precision.
Bounding Boxes Instance Extra Metrics
AP AP50 AP75 AP AP50 AP75 DICE MASD
FLOPS
Speed
Para
DiffRegFormer 52.1 71.1 57.7 46.8 71.6 52.8 56.6 3.87 0.55 1.324 140
semantic method 44.6 64.8 46.7 40.3 65.3 44.6 45.4 7.86 1.24 0.760 32.8
Table 9: Evaluation results per object class. AP denotes average precision.
Bounding Boxes - glo Instance - glo Extra Metrics
AP AP50 AP70 AP AP50 AP70 DICE MASD
DiffRegFormer 80.9 94.2 89.8 74.3 94.2 80.9 78.8 1.21
semantic method 75.6 82.4 81.2 68.3 84.3 70.1 70.6 8.68
Bounding Boxes - art Instance - art Extra Metrics
AP AP50 AP75 AP AP50 AP75 DICE MASD
DiffRegFormer 25.7 46.9 25.8 23.4 48.8 20.2 30.2 25.94
semantic method 10.3 25.8 13.6 11.5 24.9 10.3 16.3 49.14
Bounding Boxes - tub Instance - tub Extra Metrics
AP AP50 AP75 AP AP50 AP75 DICE MASD
DiffRegFormer 58.5 80.2 67.9 55.8 80.6 62.3 64.8 0.89
semantic method 43.6 73.5 54.6 46.9 75.2 52.4 50.5 2.16
Fig. 16: A visual comparison between DiffRegFormer and the semantic method. The upper row represents a zoom of the upper-box in the middle row. The lower row represents a zoom of the lower-box in the middle row. The quality of the result w.r.t. the ground truth labels can be convincingly assessed in favor of DiffRegFormer.
4.3 Complexity Assessment of Model Training